given a triangle with perimeter of 60cm and the second side is 4 more than the first and the third side is half the second. Find the length of each of the sides
Answer by atif.muhammad(135)

First side = x
Second side is 4 more than the first = (x+4)
Third side is half the second = (x+4)/2
Total length = 60
First side + second side + third side = total length
x + (x+4) + (x+4)/2 = 60
2x+4 + (x+4)/2 = 60
4x + 8 + x +4 = 120
5x + 12 = 120
5x = 108
x = 21.6 ( I think...this is a mental calculation so forgive me if i'm wrong.
Test : x + (x+4) + (x+4)/2 = 21.6 + 25.6 + 12.8 = 47.2 + 12.8 = 60, yes, it seems to be the right answer.